The information I am given is
ACT composite Scores for NVJ(a school) are normaly distrubuted with:
Population mean= 23.9422 and a population standard deviation= 4.3977
Sample Data:
26 526 28 22 31 22 26 22 28 28 28 17 16 21 29 18 35 25 25 28
Find:
Mean Median Mode Midrange Varance Standard Deviation P80 D3
i have found Mean Median Mode and Midrange but am having trouble solving for Varance Standard Deviation P80 and D3.
Can you please help me?

Answer by stanbon(75887)

Population mean= 23.9422 and a population standard deviation= 4.3977
Sample Data:
26 26 28 22 31 22 26 22 28 28 28 17 16 21 29 18 35 25 25 28
Find:
Mean Median Mode Midrange Varance Standard Deviation P80 D3
i have found Mean Median Mode and Midrange but am having trouble solving for V
Standard Deviation = 0.3233...
Variance = 0,3233^2 = 0.1045
P80 = ?
Locator : 0.8*20 = 16th number ; that data element is 28
D3= = P30
Locator : 0.3*20 = 6th number ; tat data element is 22
Can you please help me?
================================================
Cheers,
Stan H.